OpenStudy (anonymous):

8x-8 = 7x+4 x-5 = 3y-6 Solve for X? Please!

OpenStudy (anonymous):

8x - 8 = 7x + 4 Subtract 7x from both sides. x - 8 = 4 Add 8. x = 12

OpenStudy (anonymous):

Considering there is one variable y the problem might be a simultaneous equation. Which in that case your answer may be incomplete.

OpenStudy (anonymous):

solve for x and y!

OpenStudy (anonymous):

12 - 5 = 3y - 6 ( I plugged 12 which is x in) 7 = 3y - 6 13 = 3y y = 13/3

OpenStudy (anonymous):

oh gosh! u are awesome! thanks!

OpenStudy (anonymous):

welcome :)